The table shows the number of points you can score in a carnival game. Does it represent a probability distribution? explain. If it does represent a probability distribution, find the mean, variance, and standard deviation.
Points: 0, 1,2,3,4,5
Probability: .36,.21,.16,.13,.08,.06

Answer by Boreal(15235)

Discrete probabilities add to 1. Yes.
Mean is
0*0.36=0
1*0.21=0.21
2*0.16=0.32
3*0.13=0.39
4*(0.08)=0.32
5*0.06=0.30
That equals 1.54 points
variance is expected value of the squared deviation from mean
1.54^2=2.37
0.54^2=0.29
0.46^2=0.22
1.46^2=2.13
2.46^2=6.05
3.46^2=11.97
Sum is 23.03
Divide by 5=4.606
Variance
==============
SD=square root of variance=2.15